OpenStudy (anonymous):

Solve the following D.E., \[\frac{dy}{1+y^2}=\frac{dx}{1+x^2}\], obtaining the result in algebraic form

OpenStudy (anonymous):

the solution given by the text is \[y=x+C(1+xy)\]

OpenStudy (lgbasallote):

i believe you should cross multiply first \[(1 + x^2)dy - (1+y^2)dx = 0\]

myininaya (myininaya):

You just integrate both sides of that equation you have @ebbflo

myininaya (myininaya):

You already have the variables separated

OpenStudy (unklerhaukus):

\[\int \frac{1}{1+z^2}\text dz=\arctan z+c\]

myininaya (myininaya):

Is your trouble writing it as an algebraic expression?

OpenStudy (anonymous):

I see all of your points and have tried those methods but do not obtain the given answer....

myininaya (myininaya):

If so.. Don't forget you only need one constant (and just put it on one side of your equation) Remember the inverse of tan inverse which is tan lol also remember the formula tan(a+b)=(tan(a)+tan(b))/(1-tan(a)*tan(b)) Then remember that tan(constant) is still a constant)

myininaya (myininaya):

Remember @ebbflo it does say write as an algebraic expression.

myininaya (myininaya):

You get trigonometric expression right after integration

myininaya (myininaya):

You want to transform that to an algebraic one using the hint I just gave you

myininaya (myininaya):

Also recall tan(arctan(p))=p

myininaya (myininaya):

You will use that too

OpenStudy (anonymous):

thanks, I understand all the points you have made

myininaya (myininaya):

What about the solution? Have you got the desired solution?

myininaya (myininaya):

Or do you need more help?

OpenStudy (anonymous):

no, got it, i had the solution, just don't quite understand "why" the book's solution is written in the form it is... the first copyright is 1943, maybe its a style thing...

OpenStudy (anonymous):

@myininaya , you can use Latex so that your math text is more clear

myininaya (myininaya):

That one formula is a trig formula The expansion for tan(a+b) let me know if you need anything else

OpenStudy (anonymous):

as i said, I had an equivalent solution, I was just curious as to why the book chose to write the solution in the form it did...

myininaya (myininaya):

Was your form algebraic?

myininaya (myininaya):

Can you write what you have?

OpenStudy (anonymous):

I have the solution given, I think the book just wanted a form with one arbitrary constant, i generally don't like my solutions to have "y" on the LHS and RHS when it is not absolutely necessary...it was really ore of a "style" question, sorry I should have specified

myininaya (myininaya):

If the equation can be written in the form the book gives, then your answer is right I got the same answer your book got I wrote it in a different form but it is still correct

OpenStudy (anonymous):

yes I understand, thank you...as I said I was questioning the style, and now that I think of it it was probably that the books form only take a single line of text whereas mine did not...;)

OpenStudy (unklerhaukus):

differential equations will often have \(y\) on both sides of the solution, but that is alright. we were only trying to get rid of derivatives replace them with constants

OpenStudy (anonymous):

@unkle agreed, but it is my preference not to do so when not necessary

OpenStudy (anonymous):

I should have been more clear, I was not so much looking for "help" with the problem, but a different perspective on how another might write the solution

OpenStudy (unklerhaukus):

these are some solution to differential equations ,

OpenStudy (anonymous):

in order to make a decision as to what a student may find more clear

OpenStudy (unklerhaukus):

your solution must include as many arbitrary constants as the order of the differential equation

OpenStudy (anonymous):

exactly uncle, in those solution you provided, the "y" cannot be written explicitly in terms of "x", at least not in a single expression

OpenStudy (unklerhaukus):

well, it can but i just looks awful , these are 'neater'

OpenStudy (unklerhaukus):

looking at (e)

OpenStudy (anonymous):

yeah that one can but what about (c)?

OpenStudy (anonymous):

(e) more or less is already, not much more to do on that one...

OpenStudy (unklerhaukus):

i think it is clearer to write (c) like \[x = y^2(\ln y + c)\] i dont know why they chose the form they did

OpenStudy (unklerhaukus):

implicit solutions are fine for DE's